# On the application of Lorentz-Finsler geometry to model wave propagation

### v1

@@ 1 Introduction
Finsler geometry has proven to be a very powerful tool to model, in a very natural way, physical problems that exhibit an intrinsic anisotropy. The recent growing interest in the study of these models began with Zermelo’s navigation problem, first proposed in [29], but then revisited and solved in [4, 25] from a geometric point of view. In this problem, one seeks the fastest trajectory between two points for a moving object in the presence of a current. In the time-independent case, it turns out that the set of maximum velocity vectors of the object (taking into account the current) gives rise to a Finsler metric $F$—essentially, a smooth distribution of non-symmetric norms—that measures the traveltime: the $F$-length of a curve coincides with the time spent by the moving object following this path at maximum speed. The solution to Zermelo’s problem, then, can be characterized as a pregeodesic curve of $F$ (i.e. a critical point of the traveltime). [^1] A complete study of this problem via Finsler geometry can be found in [6].
This way of interpreting Zermelo’s problem provides an interesting connection with the classical Fermat’s principle, which states that the path taken by a wave ray between two points is precisely the one that makes critical the traveltime. So, if we consider a classical wave that spreads in an anisotropic medium, each individual wave trajectory can also be interpreted as a pregeodesic of a certain Finsler metric. This has led to several applications to real-world physical phenomena—especially when there is an anisotropy involved—, such as the modeling of sound waves (see [10, 11]), seismic waves (see [2, 28]) or wildfire propagation (see [8, 21]).
However, one last step to complete these models is to allow the speed of the wave to vary over time, generating a time-dependent Finsler metric. This greatly complicates the description of the problem from a purely Finslerian viewpoint due to the fact that, now, the Finsler metric no longer measures the traveltime directly and, furthermore, its pregeodesic curves cannot be defined as usual. The way to overcome this difficulty is to consider explicitly the time dimension and define a Lorentz-Finsler metric whose lightcone (the vectors with vanishing norm) represent the infinitesimal propagation of the wave through the spacetime—in the spirit of the way general relativity models the (isotropic) propagation of light using Lorentzian geometry. Then, wave trajectories can be characterized as lightlike pregeodesics of this Lorentz-Finsler metric. This provides a unified geometric framework to model wave propagation under general time-, position- and direction-dependent conditions, first developed in [14] and then applied specifically to wildfires in [15, 16].
@@ 2.1 Distance
The first thing we need to specify is the space over which the wave spreads. This space will be assumed to be a smooth manifold of dimension $n$. Since we are interested on the applications of the model and explicit computations will be needed, we will work in coordinates in $N\subset\mathds{R}^{n}$. Namely, a point in the space will be written as $x=(x^{1},\ldots,x^{n})\in N\subset\mathds{R}^{n}$ and a tangent vector as $v=(v^{1},\ldots,v^{n})\in T_{x}N\equiv\mathds{R}^{n}$. In real-world applications, this space $N$ is always endowed with a Riemannian metric $h$ (i.e. a scalar product $h_{x}(\cdot,\cdot)$ on each tangent space $T_{x}N$), which measures actual distances on $N$. This means that if $\sigma:[a,b]\rightarrow N$ is a curve on $N$, the distance we travel when following $\sigma$ is
@@
where $||v||_{h_{x}}\coloneqq\sqrt{h_{x}(v,v)}$ denotes the norm defined by $h$. This computation is independent of the parametrization of $\sigma$.
@@
    The wave spreads over a surface (in particular, a graph) of $\mathds{R}^{3}$ given by a height function $z:N\subset\mathds{R}^{2}\rightarrow\mathds{R}^{3}$. Namely, each coordinate $x=(x^{1},x^{2})\in N$ (the projection on the horizontal plane) univocally corresponds to a point $(x^{1},x^{2},z(x^{1},x^{2}))\in\mathds{R}^{3}$ of the actual surface where the propagation is taking place (see, e.g. [15, 21]). In this case, the distance is measured by the Riemmanian metric induced on the surface from the natural Euclidean metric on $\mathds{R}^{3}$:
@@
If we consider the set $\mathcal{N}_{x,y}$ of all the (piecewise smooth) curves $\sigma:[a,b]\rightarrow N$ going from $\sigma(a)=x$ to $\sigma(b)=y$, the one that minimizes the distance between $x$ and $y$ must be a critical point (i.e. it satisfies the usual Euler-Lagrange equations) of the $h$-length functional $L_{h}$, given by (1). Individual wave rays, however, do not minimize (or more generally, make critical) the distance, but the traveltime, according to the classical Fermat’s principle:
@@ 2.2 Speed profile
The initial hypotheses assume that we know the speed profile of the wave, which can be, in the most general case, rheonomic (time-dependent), inhomogeneous (position-dependent) and anisotropic (direction-dependent). This means that there exists a positive function $V(t,x,v)$ representing the speed of the wave at each (absolute) time $t\in\mathds{R}$, [^2] Here we will not consider relativistic effects, so $t$ is absolute in the sense that it is the time measured by every observer. each point $x\in N$ and each direction $v\in T_{x}N$. Obviously, this speed should not depend on the length of $v$, only on the direction it points, so $V$ must be positive homogeneous of degree 00 on the direction: [^3] It will be assumed that the wave can spread in all directions. The case of a “strong wind” effect that makes the wave unable to propagate in some directions is studied in detail in [14, § 6].
@@
For example, if $n=2$ we can express the speed profile as $V(t,x,\theta)$, where $\theta$ denotes the angle in the usual polar coordinates (see Example 2(i) below).
@@
$V(t(s),\sigma(s),\dot{\sigma}(s))=\left\lvert\left\lvert\frac{d(\sigma\circ s)}{dt}(t(s))\right\rvert\right\rvert_{h_{\sigma(s)}}=\dot{s}(t(s))||\dot{\sigma}(s)||_{h_{\sigma(s)}},$
@@ 2.3 Finsler metrics
  • $F$ is smooth away from the zero vector, so long $V$ is smooth.
@@
    which is a hypersurface of $T_{x}N$ diffeomorphic to a sphere and enclosing the zero vector (e.g. when $n=2$ the indicatrix is a simple closed curve enclosing the origin).
  • We define the fundamental tensor of $F_{t,x}$ in the direction $v\in T_{x}N\setminus\{0\}$, denoted $g_{v}^{F_{t,x}}$, as the symmetric bilinear form given by the Hessian of $\frac{1}{2}F^{2}_{t,x}$ at $v$: [^4] This is the way one obtains a metric out of a norm. For instance, the Hessian of $\frac{1}{2}||\cdot||_{h}^{2}$ provides the Riemannian metric $h$.
@@
These two objects are related by the following result (see, e.g. [17, Proposition 2.3(v)]).
@@
If this holds for every $(t,x)\in\mathds{R}\times N$, then $F$ is said to be a (time-dependent) Finsler metric on $N$ [^5] See [17] for the formal definition and properties of a Finsler metric.
When $F$ is a Finsler metric, it satisfies the strict triangle inequality, which is the last basic property usually required to norms: [^6] In fact, this inequality holds so long $\Sigma_{t,x}$ is strictly convex (see [17, Proposition 2.3(iv)]).
@@
with equality if and only if $v=\lambda u$ for some $\lambda>0$. This way, fixing $(t,x)\in\mathds{R}\times N$, we can think of $F_{t,x}$ as a non-symmetric norm (in general $F_{t,x}(v)\not=F_{t,x}(-v)$), the fundamental tensor $g_{v}^{F_{t,x}}$ as its associated scalar product for each direction $v$, and the indicatrix $\Sigma_{t,x}$ represents the infinitesimal propagation of the wave: $\Sigma_{t,x}$ is the wavefront one would obtain after one time unit if the wave departed from $(t,x)$ under constant conditions (i.e. if $F$ were time- and position-independent).
@@
    $F_{t,x}(v)=\frac{\sqrt{(v^{1})^{2}+(v^{2})^{2}}}{a(t,x)(1-\varepsilon(t,x)^{2})}\left[1-\varepsilon(t,x)\cos\left(\arctan\left(\frac{v^{2}}{v^{1}}\right)-\phi(t,x)\right)\right],$
@@
    which is a Matsumoto metric (see, e.g. [26]), used to model wave propagation that takes place on a surface and is affected by the slope. Choosing the $+$ sign means that the propagation is faster upwards (as is the case for wildfires; see [15, § 3.1]), whereas the -- sign means that it is faster downwards.
Geodesic curves will play a key role in this model, as we will see later. When $F$ is time-independent, we say that it is a proper Finsler metric on $N$ and, similar to the case with Riemannian metrics, we can define the geodesics of $F$ as the critical points of the $F$-energy functional (see [5])
[…]
However, $E_{F}[\sigma]$ is preserved only under positive affine reparametrizations of $\sigma$, [^8] Specifically, $\sigma(s)$ is a geodesic of $F$ if and only if so is $\sigma\circ s(r)$, with $s(r)=\lambda r+c$ for some $\lambda>0$ and $c\in\mathds{R}$. i.e. geodesics require a specific parametrization—essentially, they represent trajectories without acceleration—, whereas here we will be simply interested on the path they follow. In order to remove this dependence on the parametrization, we define the pregeodesics of $F$ as the curves that can be reparametrized as geodesics (keeping the orientation). It turns out that pregeodesics are the critical points of the $F$-length functional (see [3, § 5])
[…]
which is indeed independent of reparametrizations of $\sigma$, thanks to the fact that $F$ is positive homogeneous of degree 1 and time-independent. In the general time-dependent case, however, $E_{F}[\sigma]$ is not preserved even under affine reparametrizations and $L_{F}[\sigma]$ also becomes dependent on the parametrization, so (pre-)geodesics cannot be defined as usual and we need to generalize this notion.
@@ 2.4 Lorentz-Finsler metrics
To deal with the time dependence and solve the issue of defining geodesic curves in this case, it is more convenient to treat the time explicitly as an additional dimension and consider the spacetime $M\coloneqq\mathds{R}\times N$, where the natural projection (time coordinate) $t:M\rightarrow\mathds{R}$ measures the absolute time. This way, points in the spacetime will be expressed as $p=(t,x)\in M$, vectors as $\hat{v}=(v^{0},v)\in T_{p}M\equiv\mathds{R}\times T_{x}N$ and curves $\gamma:[a,b]\rightarrow M$ as $\gamma(s)=(t(s),\sigma(s))$. The idea then is to mimick the way general relativity models the propagation of light, constructing a metric with Lorentzian signature (+,,,)(+,-$\ldots$,-)( + , - , … , - ) where the curves with vanishing norm represent the trajectories of individual wave rays.
@@
Observe that, for any $v\in T_{x}N\setminus\{0\}$, there exists a unique $v^{0}>0$ such that $\hat{v}=(v^{0},v)$ is lightlike. In fact, $v\in\Sigma_{t,x}$ if and only if $(1,v)\in\mathcal{C}_{t,x}$. Analogously, any curve $\sigma(s)$ on $N$ can be lifted to a unique lightlike curve $\gamma(s)=(t(s),\sigma(s))$ on $M$.
@@
  • $G$ is positive homogeneous of degree 2: $G_{p}(\lambda\hat{v})=\lambda^{2}G_{p}(\hat{v})$, for all $\lambda>0$. The reason to use a 2-homogeneous function instead of a 1-homogeneous one such as $F$ (and thus, the reason why we squared equation (7) to obtain $G$) is because 1-homogeneous functions fail to be smooth on lightlike vectors.
@@
    $\begin{split}g^{G_{p}}_{\hat{v}}(\hat{u},\hat{w})&\coloneqq\frac{1}{2}\textup{Hess}(G_{p}(\hat{v}))(\hat{u},\hat{w})=\frac{1}{2}\left.\frac{\partial^{2}}{\partial\delta\partial\eta}G_{p}(\hat{v}+\delta\hat{u}+\eta\hat{w})\right\rvert_{\delta=\eta=0}\\ &=u^{0}w^{0}-g^{F_{p}}_{v}(u,w).\end{split}$ (10)
@@
    $\begin{split}g^{G_{p}}_{\hat{v}}(\hat{u},\hat{w})&=\frac{1}{2}\begin{pmatrix}u^{0}&\cdots&u^{n}\end{pmatrix}\begin{pmatrix}\frac{\partial^{2}G_{p}}{\partial(v^{0})^{2}}(\hat{v})&\cdots&\frac{\partial^{2}G_{p}}{\partial v^{0}\partial v^{n}}(\hat{v})\\ \vdots&\ddots&\vdots\\ \frac{\partial^{2}G_{p}}{\partial v^{n}\partial v^{0}}(\hat{v})&\cdots&\frac{\partial^{2}G_{p}}{\partial(v^{n})^{2}}(\hat{v})=\end{pmatrix}\begin{pmatrix}w^{0}\\ \vdots\\ w^{n}\end{pmatrix}\\ &=\frac{1}{2}\begin{pmatrix}u^{0}&\cdots&u^{n}\end{pmatrix}\begin{pmatrix}1&0&\ldots&0\\ 0&-\frac{\partial^{2}F^{2}_{p}}{\partial(v^{1})^{2}}(v)&\cdots&-\frac{\partial^{2}F^{2}_{p}}{\partial v^{1}\partial v^{n}}(v)\\ \vdots&\vdots&\ddots&\vdots\\ 0&-\frac{\partial^{2}F^{2}_{p}}{\partial v^{n}\partial v^{1}}(v)&\cdots&-\frac{\partial^{2}F^{2}_{p}}{\partial(v^{n})^{2}}(v)=\end{pmatrix}\begin{pmatrix}w^{0}\\ \vdots\\ w^{n}\end{pmatrix}.\end{split}$ (11)
    Note that $g^{G_{p}}_{\hat{v}}(\hat{v},\hat{v})=G_{p}(\hat{v})$ and $g^{G_{p}}_{\hat{v}}(\cdot,\cdot)$ is a Lorentzian scalar product with signature (+,,,)(+,-$\ldots$,-)( + , - , … , - ).
Now we can define the geodesics of $G$ as the critical points of the $G$-energy functional (see [19])
@@
which is preserved under positive affine reparametrizations of $\gamma$, even if $F$ is time-dependent. Note that we cannot define a length functional in this case, as $\sqrt{G}$ is not smooth on lightlike vectors, but we can still define the pregeodesics of $G$ as the curves that can be reparametrized as geodesics (keeping the orientation). Of course, when $F$ is time-independent, then $\gamma(s)=(t(s),\sigma(s))$ is a pregeodesic of $G$ if and only if $\sigma(s)$ is a pregeodesic of $F$ (see [7, Proposition B.1]). This way, we can see the pregeodesics of $G$ (or more specifically, their projection on $N$) as a more general version of those of $F$.

### v2

# Enrique Pendás-Recondo

@@ 1 Introduction
Finsler geometry has proven to be a very powerful tool to model, in a very natural way, physical problems that exhibit an intrinsic anisotropy. The recent growing interest in the study of these models began with Zermelo’s navigation problem, first proposed in [29], but then revisited and solved in [4, 25] from a geometric point of view. In this problem, one seeks the fastest trajectory between two points for a moving object in the presence of a current. In the time-independent case, it turns out that the set of maximum velocity vectors of the object (taking into account the current) gives rise to a Finsler metric FF—essentially, a smooth distribution of non-symmetric norms—that measures the traveltime: the $F$-length of a curve coincides with the time spent by the moving object following this path at maximum speed. The solution to Zermelo’s problem, then, can be characterized as a pregeodesic curve of $F$ (i.e., a critical point of the traveltime). [^1] A complete study of this problem via Finsler geometry can be found in [6].
This way of interpreting Zermelo’s problem provides an interesting connection with the classical Fermat’s principle, which states that the path taken by a wave ray between two points is precisely the one that makes critical the traveltime. So, if we consider a classical wave that spreads in an anisotropic medium, each individual wave trajectory can also be interpreted as a pregeodesic of a certain Finsler metric. This has led to several applications to real-world physical phenomena—especially when there is an anisotropy involved—such as the modeling of sound waves (see [10, 11]), seismic waves (see [2, 28]) or wildfire propagation (see [8, 21]).
However, one last step to complete these models is to allow the speed of the wave to vary over time, generating a time-dependent Finsler metric. This greatly complicates the description of the problem from a purely Finslerian viewpoint due to the fact that, now, pregeodesic curves on the space cannot be defined as usual. The way to overcome this difficulty is to add explicitly the time dimension to the space and define a Lorentz-Finsler metric whose lightcone (the vectors with vanishing norm) represents the infinitesimal propagation of the wave through the resulting spacetime—in the spirit of the way general relativity models the (isotropic) propagation of light using Lorentzian geometry. Then, wave trajectories can be characterized as lightlike pregeodesics of this Lorentz-Finsler metric. This provides a unified geometric framework to model wave propagation under general time-, position- and direction-dependent conditions, first developed in [14] and then applied specifically to wildfires in [15, 16].
@@ 2.1 Distance
The first thing we need to specify is the space over which the wave spreads. This space will be assumed to be a smooth manifold of dimension $n$. Since we are interested on the applications of the model and explicit computations will be needed, we will work in coordinates in $N\subset\mathds{R}^{n}$. Namely, a point in the space will be written as $x=(x^{1},\ldots,x^{n})\in N\subset\mathds{R}^{n}$ and a tangent vector as $v=(v^{1},\ldots,v^{n})\in T_{x}N\equiv\mathds{R}^{n}$. In real-world applications, this space $N$ is always endowed with a Riemannian metric $h$ (i.e., a scalar product $h_{x}(\cdot,\cdot)$ on each tangent space $T_{x}N$), which measures actual distances on $N$. This means that if $\sigma:[a,b]\rightarrow N$ is a curve on $N$, the distance we travel when following $\sigma$ is
@@
where $v||_{h_{x}}\coloneqq\sqrt{h_{x}(v,v)}$ denotes the norm defined by $h$. This computation is independent of the parametrization of $\sigma$.
@@
    The wave spreads over a surface (in particular, a graph) of $\mathds{R}^{3}$ given by a height function $z:N\subset\mathds{R}^{2}\rightarrow\mathds{R}^{3}$. Namely, each coordinate $x=(x^{1},x^{2})\in N$ (the projection on the horizontal plane) univocally corresponds to a point $(x^{1},x^{2},z(x^{1},x^{2}))\in\mathds{R}^{3}$ of the actual surface where the propagation is taking place (see e.g. [15, 21]). In this case, the distance is measured by the Riemmanian metric induced on the surface from the natural Euclidean metric on $\mathds{R}^{3}$:
@@
If we consider the set $\mathcal{N}_{x,y}$ of all (piecewise smooth) curves $\sigma:[a,b]\rightarrow N$ going from $\sigma(a)=x$ to $\sigma(b)=y$, the one that minimizes the distance between $x$ and $y$ must be a critical point (i.e., it satisfies the usual Euler-Lagrange equations) of the $h$-length functional $L_{h}$, given by (1). Individual wave rays, however, do not minimize (or more generally, make critical) the distance, but the traveltime, according to the classical Fermat’s principle:
@@ 2.2 Speed profile
The initial hypotheses assume that we know the speed profile of the wave, which can be, in the most general case, rheonomic (time-dependent), inhomogeneous (position-dependent) and anisotropic (direction-dependent). This means that there exists a positive function $V(t,x,v)$ representing the speed of the wave at each (absolute) time $t\in\mathds{R}$, [^2] Here we will not consider relativistic effects, so $t$ is absolute in the sense that it is the time measured by every observer. each point $x\in N$ and each direction $v\in T_{x}N$. Obviously, this speed should not depend on the length of $v$, only on the direction it points, so $V$ must be positive homogeneous of degree 0 on the direction: [^3] It will be assumed that the wave can spread in all directions. The case of a “strong wind” effect that makes the wave unable to propagate in some directions is studied in detail in [14, § 6].
@@
For example, if $n=2$ we can express the speed profile as $V(t,x,\theta)$, where $\theta\in[0,2\pi)$ denotes the angle in the usual polar coordinates (see Example 2(i) below).
@@
[…]
@@ 2.3 Finsler metrics
  • $F$ is smooth away from the zero vector, as long as $V$ is smooth.
@@
    which is a hypersurface of $T_{x}N$ diffeomorphic to a sphere and enclosing the zero vector (e.g., when $n=2$, the indicatrix is a simple closed curve enclosing the origin).
  • We define the fundamental tensor of $F_{t,x}$ in the direction $v\in T_{x}N\setminus\{0\}$, denoted $g_{v}^{F_{t,x}}$, as the symmetric bilinear form given by the Hessian of $\frac{1}{2}F^{2}_{t,x}$ at vv: [^4] This is the way one obtains a metric out of a norm. For instance, the Hessian of $\frac{1}{2}||\cdot||_{h}^{2}$ provides the Riemannian metric $h$.
@@
These two objects are related to each other by the following result (see e.g. [17, Proposition 2.3(v)]).
@@
If this holds for every $(t,x)\in\mathds{R}\times N$, then $F$ is said to be a (time-dependent) Finsler metric on NN. [^5] See [17] for the formal definition and properties of a Finsler metric.
When $F$ is a Finsler metric, it satisfies the strict triangle inequality, which is the last basic property usually required to norms: [^6] In fact, this inequality holds as long as $\Sigma_{t,x}$ is strictly convex (see [17, Proposition 2.3(iv)]).
@@
with equality if and only if $v=\lambda u$ for some $\lambda>0$. This way, fixing $(t,x)\in\mathds{R}\times N$, we can think of $F_{t,x}$ as a non-symmetric norm (in general $F_{t,x}(v)\not=F_{t,x}(-v)$), the fundamental tensor $g_{v}^{F_{t,x}}$ as its associated scalar product for each direction $v$, and the indicatrix $\Sigma_{t,x}$ represents the infinitesimal propagation of the wave: $\Sigma_{t,x}$ is the wavefront one would obtain after one time unit if the wave departed from $(t,x)$ under constant conditions (i.e., if $F$ were time- and position-independent).
@@
    $F_{t,x}(v)=\frac{\sqrt{(v^{1})^{2}+(v^{2})^{2}}}{a(t,x)(1-\varepsilon(t,x)^{2})}\left[1-\varepsilon(t,x)\cos\left(\operatorname{atan2}(v^{2},v^{1})-\phi(t,x)\right)\right],$
@@
    which is a Matsumoto metric (see e.g. [26]), used to model wave propagation that takes place on a surface and is affected by the slope. Choosing the ++ sign means that the propagation is faster upwards (as is the case for wildfires; see [15, § 3.1]), whereas the - sign means that it is faster downwards.
Geodesic curves will play a key role in this model, as we will see later. When $F$ is time-independent, we say that it is a proper Finsler metric on $N$ and, similar to the case with Riemannian metrics, we can define the geodesics of $F$ as the critical points of the $F$-energy functional
$E_{F}[\sigma]\coloneqq\int_{a}^{b}F_{\sigma(s)}(\dot{\sigma}(s))^{2}ds,$
among all (piecewise smooth) curves $\sigma:[a,b]\rightarrow N$ with fixed endpoints (see e.g. [5]). However, these critical points are preserved only under positive affine reparametrizations of $\sigma$, [^8] Specifically, $\sigma(s)$ is a geodesic of $F$ if and only if so is $\sigma\circ s(r)$, with $s(r)=\lambda r+c$ for some $\lambda>0$ and $c\in\mathds{R}$. i.e. geodesics require a specific parametrization—essentially, they represent trajectories without acceleration—whereas here we are simply interested on the path they follow. In order to remove this dependence on the parametrization, we define the pregeodesics of $F$ as the curves that can be reparametrized as geodesics (keeping the orientation). It turns out that pregeodesics are the critical points of the $F$-length functional (see [3, § 5])
$L_{F}[\sigma]\coloneqq\int_{a}^{b}F_{\sigma(s)}(\dot{\sigma}(s))ds,$ (6)
which is indeed independent of reparametrizations of $\sigma$, thanks to the fact that $F$ is positive homogeneous of degree 1.
In the general time-dependent case, however, $E_{F}$ and $L_{F}$ not only depend on $\sigma(s)$ but also on the external time function $t(s)$. Therefore, (pre-)geodesic curves cannot be defined as usual—the same curve $\sigma$ on $N$ can have different $F$-energy and $F$-length in different contexts—and we need to generalize this notion.
@@ 2.4 Lorentz-Finsler metrics
To deal with the time dependence and solve the issue of defining geodesic curves in this case, time must be included as an intrinsic element of the manifold, rather than as an external function. This can be achieved by explicitly treating time as an additional dimension and considering the spacetime $M\coloneqq\mathds{R}\times N$, where the natural projection (time coordinate) $t:M\rightarrow\mathds{R}$ measures the absolute time. This way, points in the spacetime will be expressed as $p=(t,x)\in M$, vectors as $\hat{v}=(v^{0},v)\in T_{p}M\equiv\mathds{R}\times T_{x}N$ and curves $\gamma:[a,b]\rightarrow M$ as $\gamma(s)=(t(s),\sigma(s))$. The idea then is to mimic the way general relativity models the propagation of light, constructing a metric with Lorentzian signature $(+,-,\ldots,-)$ where curves with vanishing norm represent trajectories of individual wave rays.
@@
Observe that, for any $v\in T_{x}N\setminus\{0\}$ and $t\in\mathds{R}$, there exists a unique $v^{0}>0$ such that $\hat{v}=(v^{0},v)\in\mathcal{C}_{t,x}$. In fact, $v\in\Sigma_{t,x}$ if and only if $(1,v)\in\mathcal{C}_{t,x}$.
@@
[…]
@@
    $\begin{split}g^{G_{p}}_{\hat{v}}(\hat{u},\hat{w})&\coloneqq\frac{1}{2}\textup{Hess}(G_{p}(\hat{v}))(\hat{u},\hat{w})=\frac{1}{2}\left.\frac{\partial^{2}}{\partial\delta\partial\eta}G_{p}(\hat{v}+\delta\hat{u}+\eta\hat{w})\right\rvert_{\delta=\eta=0}\\ &=u^{0}w^{0}-g^{F_{p}}_{v}(u,w).\end{split}$
@@
[…]
Now we can define the geodesics of $G$ as the critical points of the $G$-energy functional
$E_{G}[\gamma]\coloneqq\int_{a}^{b}G_{\gamma(s)}(\dot{\gamma}(s))ds,$ (11)
among all (piecewise smooth) curves $\gamma:[a,b]\rightarrow M$ with fixed endpoints (see e.g. [19]). These critical points are preserved under positive affine reparametrizations of $\gamma$, even if $F$ is time-dependent. Although we cannot characterize in general pregeodesic curves as critical points of a length functional, as before, since $\sqrt{G}$ can be non-smooth on lightlike vectors, we can still define the pregeodesics of $G$ as the curves that can be parametrized as geodesics (keeping the orientation). Of course, when $F$ is time-independent, then $\gamma(s)=(t(s),\sigma(s))$ is a pregeodesic of $G$ if and only if $\sigma(s)$ is a pregeodesic of $F$ (see [7, Proposition B.1]). This way, we can see pregeodesics of $G$ (or more specifically, their projection on $N$) as a generalization of those of $F$ in the time-dependent case.